# RDEIC: Accelerating Diffusion-Based Extreme Image Compression with Relay Residual Diffusion

### v2

@@ IV-A2 Implementation details
We use Stable Diffusion 2.1-base [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the specific implementation of stable diffusion. Throughout all our experiments, the weights of stable diffusion remain frozen. Similar to DiffEIC [16], the control module in our RDEIC has the same encoder and middle block architecture as stable diffusion and reduces the channel number to 20% of the original. The variance sequence $\{\beta_{t}\}_{t=1}^{T}$ used for adding noise is identical to that in Stable Diffusion. The number $N$ of denoising steps is set to 300. The size of codebook is set to 16384.
For training, we use the Adam [48] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for a total of 300K iterations. To achieve different compression ratios, we train five models with $\lambda_{r}$ selected from {2, 1, 0.5, 0.25, 0.1}. The batch size is set to 4. As described in Section III-C, the training process is divided into two stages. 1) Independent training. During this stage, the initial learning rate is set to 1×$\times 10^{-4}$ and images are randomly cropped to 512$\times$512 patches. We first train the proposed RDEIC with $\lambda_{r}=2$ for 100K iterations. The learning rate is then reduced to 2×$\times 10^{-5}$ and the model is trained with target $\lambda_{r}$ for another 100K iterations. 2) Fixed-step fine-tuning. In this stage, the learning rate is set to 2×$\times 10^{-5}$ and images are randomly cropped to 256$\times$256 patches. We fine-tune the model through the entire reconstruction process for 100K iterations. When $\lambda_{r}\in\{2,1\}$, the fixed number $L$ is set to 2, otherwise, it is 5. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
@@ IV-B Comparisons With State-of-the-art Methods
We compare the proposed RDEIC with several representative extreme image compression methods, including the traditional standards: BPG [2] and VVC [3]; VAE-based method: ELIC [20]; GANs-based methods: HiFiC [9], MS-ILLM [10], and VQIR [29]; and diffusion-based methods: Text+Sketch [14], PerCo [15],and DiffEIC [16].
Original VVC MS-ILLM Text+Sketch PerCo DiffEIC RDEIC (Ours)
[…]
Figure 6: Visual comparisons of our method to baselines on the CLIC2020 dataset. Compared to other methods, our method produces more realistic and faithful reconstructions.
TABLE I: Encoding and decoding time (in seconds) on Kodak dataset. Decoding time is divided into the time spent in the denoising stage and the time spent in the remaining parts. The testing platform is RTX4090.
Types Methods Denoising Step Encoding Time Decoding time
@@
VAE-based ELIC 0.056 $\pm$ 0.006   – 0.081 $\pm$ 0.011
GAN-based HiFiC 0.038 $\pm$ 0.004   – 0.059 $\pm$ 0.004
MS-ILLM 0.038 $\pm$ 0.004   – 0.059 $\pm$ 0.004
VQIR 0.050 $\pm$ 0.003   – 0.179 $\pm$ 0.005
Diffusion-based Text+Sketch 25 62.045 $\pm$ 0.516 8.483 $\pm$ 0.344 4.030 $\pm$ 0.469
DiffEIC 50 0.128 $\pm$ 0.005 4.342 $\pm$ 0.013 0.228 $\pm$ 0.026
PerCo 5 0.236 $\pm$ 0.040 0.623 $\pm$ 0.003 0.186 $\pm$ 0.002
20 0.236 $\pm$ 0.040 2.495 $\pm$ 0.009 0.186 $\pm$ 0.002
RDEIC (Ours) 2 0.119 $\pm$ 0.003 0.173 $\pm$ 0.001 0.198 $\pm$ 0.003
5 0.119 $\pm$ 0.003 0.434 $\pm$ 0.002 0.198 $\pm$ 0.003
@@ IV-B1 Quantitative Comparisons
Fig. 4 shows the rate-perception curves of the proposed and compared methods across the three datasets. It can be observed that the proposed RDEIC demonstrates superior performance across different perceptual metrics compared to other methods, particularly achieving optimal results in DISTS, FID, and KID. For the distortion metrics, as shown in Fig. 5, RDEIC significantly outperforms other diffusion-based methods, underscoring its superiority in maintaining consistency. Moreover, we report the performance of the SD autoencoder in Fig. 4 and Fig. 5 (see the black horizontal line), which represents the upper bound of RDEIC’s performance. Compared to DiffEIC [16], which is also based on stable diffusion, RDEIC is significantly closer to this performance upper limit.
@@ IV-B3 Complexity Comparisons
Figure 7: Ablation studies of the proposed relay residual diffusion and fixed-step fine-tuning on the CLIC2020 dataset.
TABLE II: The impact of RRD and FSFT on performance. BD-Rate is evaluated on CLIC2020 dataset, using DiffEIC [16] as the anchor. DS denotes the number of denoising steps. 2/5 denotes that we use 2 denoising steps for the two models with larger bpp and 5 steps for the remaining models.
Methods DS BD-Rate (%)
DISTS LPIPS PSNR MS-SSIM
DiffEIC 50 0 0 0 0
Baseline 50 1.57 4.88 9.72 3.58
+RRD 2/5 12.63 -4.52 -47.24 -18.65
+RRD+FSFT 2/5 -17.94 -40.79 -61.26 -32.73
@@ V Analysis and Discussions
To provide a more comprehensive analysis of the proposed method, we conduct ablation studies, with the results presented in Fig. 7 and Table II. For the baseline, we employ the same diffusion framework as DiffEIC [16], where the denoising process starts from pure noise. As shown in Fig. 7, our baseline performs similarly to DiffEIC [16], indicating that the choice of compression module has minimal impact on overall performance.
@@ V-A Effectiveness of Relay Residual Diffusion
We first investigate the effectiveness of our proposed relay residual diffusion framework. As shown in Fig. 7 and Table II, by incorporating the proposed relay residual diffusion framework, we achieve better distortion performance and comparable perceptual performance with 2/5 denoising steps compared to the Baseline, which uses 50 denoising steps. The reason behind this is that starting from the compressed latent feature, instead of pure noise, avoids the error accumulation in the initial stage of the denoising process and provides a solid foundation for subsequent detail generation.
@@
Since the time required for the denoising stage is directly proportional to the number of denoising steps, incorporating RRD reduces the denoising time by a factor of 10$\times$ to 25$\times$ compared to the baseline, as shown in Table III.
TABLE III: The impact of RRD on speed. DS denotes the number of denoising steps. FSFT is a fine-tuning strategy that does not affect speed.
Methods DS Denoising Time Speedup
DiffEIC 50 4.342 $\pm$ 0.013 1$\times$
Baseline 50 4.349 $\pm$ 0.013 1$\times$
+RRD 5 0.434 $\pm$ 0.002 10$\times$
2 0.173 $\pm$ 0.001 25$\times$
V-B Analyze of Denoising Steps
Next, we analyze the impact of denoising steps on “Baseline+RRD” to select an appropriate value of $L$ for FSFT strategy. As shown in Fig. 8, for $\lambda_{r}\in\{2,1\}$, the number of denoising steps has minimal effect on compression performance, so that we set $L$ to 2 in this case. For $\lambda_{r}\in\{0.5,0.25,0.1\}$, increasing the denoising steps achieves better perceptual results (lower LPIPS and DISTS values), but leads to degraded fidelity (lower PSNR and MS-SSIM values). To achieve a balance between fidelity and perceptual quality, we set $L$ to 5 here.
Figure 8: The impact of denoising steps on “Baseline+RRD”.
@@ V-B Analyze of Denoising Steps
V-C Effectiveness of Fixed-step Fine-tuning
We further demonstrate the effectiveness of the proposed FSFT strategy. As shown in Fig. 7 and Table II, the FSFT strategy significantly improves reconstruction performance across all metrics, indicating that it effectively eliminates the discrepancy between the training and inference phases. Furthermore, as FSFT is a fine-tuning strategy, it does not introduce any additional computational overhead during inference.
Original $\lambda_{s}$=0.0 $\lambda_{s}$=0.6 $\lambda_{s}$=1.0 $\lambda_{s}$=1.3 $\lambda_{s}$=1.5
Refer to caption Refer to caption
Figure 10: Balancing smoothness versus sharpness. The second and fourth rows show the absolute difference between the reconstructed images and the baseline ($\lambda_{s}=0$).
V-D Smoothness-sharpness Trade-off
V-E Role of the Diffusion Mechanism
V-F Limitations
Original W/o diffusion mechanism W/o denoising process RDEIC
@@ VI Conclusion
In this paper, we propose an innovative relay residual diffusion-based method (RDEIC) for extreme image compression. Unlike most existing diffusion-based methods that start from pure noise, RDEIC takes the compressed latent features of the input image with added noise as the starting point and reconstructs the image by iteratively removing the noise and reducing the residual between the compressed latent features and the target latent features. Extensive experiments have demonstrated the superior performance of our RDEIC over existing state-of-the-art methods in terms of both reconstruction quality and computational complexity.

### v3

@@ IV-A2 Implementation details
We use Stable Diffusion 2.1-base [^1] https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the specific implementation of stable diffusion. Throughout all our experiments, the weights of stable diffusion remain frozen. Similar to DiffEIC [16], the control module in our RDEIC has the same encoder and middle block architecture as stable diffusion and reduces the channel number to 20% of the original. The variance sequence $\{\beta_{t}\}_{t=1}^{T}$ used for adding noise is identical to that in Stable Diffusion. The number $N$ of denoising steps is set to 300. The size of codebook is set to 16,384.
For training, we use the Adam [52] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for a total of 300K iterations. To achieve different compression ratios, we train five models with $\lambda_{r}$ selected from {2, 1, 0.5, 0.25, 0.1}. The batch size is set to 4. As described in Section III-C, the training process is divided into two stages. 1) Independent training. During this stage, the initial learning rate is set to 1×$\times 10^{-4}$ and images are randomly cropped to 512$\times$512 patches. We first train the proposed RDEIC with $\lambda_{r}=2$ for 100K iterations. The learning rate is then reduced to 2×$\times 10^{-5}$ and the model is trained with target $\lambda_{r}$ for another 100K iterations. 2) Fixed-step fine-tuning. In this stage, the learning rate is set to 2×$\times 10^{-5}$. To reduce the computational burden, images are randomly cropped to 256$\times$256 patches. We fine-tune the model through the entire reconstruction process for 100K iterations. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU.
Original VVC MS-ILLM Text+Sketch PerCo DiffEIC RDEIC (Ours)
[…]
Figure 7: Visual comparisons of our method to baselines on the CLIC2020 dataset. Compared to other methods, our method produces more realistic and faithful reconstructions. For our RDEIC, the first example corresponds to the result of RDEIC-5, while the second example corresponds to RDEIC-2.
@@ IV-B Comparisons With State-of-the-art Methods
We compare the proposed RDEIC with several representative extreme image compression methods, including the traditional standards: BPG [2] and VVC [3]; VAE-based method: ELIC [20]; GANs-based methods: HiFiC [9], MS-ILLM [10], and VQIR [33]; and diffusion-based methods: Text+Sketch [14], PerCo [15],and DiffEIC [16]. For our RDEIC, we present the performance curves of models fine-tuned with 2 and 5 fixed denoising steps, i.e., RDEIC-2 and RDEIC-5.
TABLE I: Encoding and decoding time (in seconds) on Kodak dataset. Decoding time is divided into the time spent in the denoising stage and the time spent in the remaining parts. BD-Rate is evaluated on the CLIC2020 dataset using DISTS as the metric, with VVC [3] serving as the anchor. The best and second best results are highlighted in bold and underline. The testing platform is RTX4090.
Types Methods Denoising Step Encoding Time Decoding time BD-Rate(%$\downarrow$
@@
VAE-based ELIC 0.056 $\pm$ 0.006   – 0.081 $\pm$ 0.011 20.06
GAN-based HiFiC 0.038 $\pm$ 0.004   – 0.059 $\pm$ 0.004 -48.11
MS-ILLM 0.038 $\pm$ 0.004   – 0.059 $\pm$ 0.004 -88.72
VQIR 0.050 $\pm$ 0.003   – 0.179 $\pm$ 0.005 -75.49
Diffusion-based Text+Sketch 25 62.045 $\pm$ 0.516 8.483 $\pm$ 0.344 4.030 $\pm$ 0.469 111.94
DiffEIC 50 0.128 $\pm$ 0.005 4.342 $\pm$ 0.013 0.228 $\pm$ 0.026 -86.86
PerCo 5 0.236 $\pm$ 0.040 0.623 $\pm$ 0.003 0.186 $\pm$ 0.002 -76.70
20 0.236 $\pm$ 0.040 2.495 $\pm$ 0.009 0.186 $\pm$ 0.002 -76.70
RDEIC (Ours) 2 0.119 $\pm$ 0.003 0.173 $\pm$ 0.001 0.198 $\pm$ 0.003 -89.44
5 0.119 $\pm$ 0.003 0.434 $\pm$ 0.002 0.198 $\pm$ 0.003 -94.27
TABLE II: The impact of Relay Residual Diffusion (RRD) and Fixed-Step Fine-Tune (FSFT) on performance. BD-Rate is evaluated on CLIC2020 dataset, using DiffEIC [16] as the anchor. DS denotes the number of denoising steps.
Methods DS BD-Rate (%) $\downarrow$ Denoising Time Speedup
DISTS LPIPS PSNR MS-SSIM
DiffEIC 50 0 0 0 0 4.342 $\pm$ 0.013 1$\times$
Baseline 50 1.57 4.88 9.72 3.58 4.349 $\pm$ 0.013 1$\times$
+RRD 5 12.43 -4.50 -47.45 -18.70 0.434 $\pm$ 0.002 10$\times$
2 36.33 13.83 -56.75 -26.21 0.173 $\pm$ 0.001 25$\times$
RDEIC (+RRD+FSFT) 5 -18.02 -40.79 -61.11 -32.73 0.434 $\pm$ 0.002 10$\times$
2 -15.08 -39.07 -61.27 -31.89 0.173 $\pm$ 0.001 25$\times$
@@ IV-B1 Quantitative Comparisons
Fig. 5 shows the rate-perception curves of the proposed and compared methods across the three datasets. It can be observed that the proposed RDEIC demonstrates superior performance across different perceptual metrics compared to other methods, particularly achieving optimal results in DISTS, FID, and KID. For the distortion metrics, as shown in Fig. 6, RDEIC significantly outperforms other diffusion-based methods, underscoring its superiority in maintaining consistency. RDEIC-2 and RDEIC-5 achieve comparable performance across most metrics. However, at the lowest bitrate setting (i.e., $\lambda_{r}=0.1$), using more denoising steps results in better perceptual quality, as reflected in the perceptual metrics. Moreover, we report the performance of the SD autoencoder in Fig. 5 and Fig. 6 (see the black horizontal line), which represents the upper bound of RDEIC’s performance, as it reflects the best achievable reconstruction quality without compression. Compared to DiffEIC [16], which is also based on stable diffusion, RDEIC is significantly closer to this performance upper limit.
@@ V Analysis and Discussions
To provide a more comprehensive analysis of the proposed method, we conduct ablation studies. For the baseline, we employ the same diffusion framework as DiffEIC [16], where the denoising process starts from pure noise. As shown in Table II, our baseline performs similarly to DiffEIC [16], indicating that the performance improvements of our method primarily stem from the proposed relay residual diffusion (RRD) framework and fixed-step fine-tuning (FSFT) strategy, while the choice of compression module has minimal impact on the overall performance.
(a) Original (b) Baseline-2 (c) +RRD-2 (d) +RRD+FSFT-2
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 8: Visual comparisons of each component in the proposed method.
Original $\lambda_{s}$=0.0 $\lambda_{s}$=0.6 $\lambda_{s}$=0.8 $\lambda_{s}$=1.0 $\lambda_{s}$=1.3 $\lambda_{s}$=1.5
[…]
Figure 9: Balancing smoothness versus sharpness. These results are produced by RDEIC-2 trained with $\lambda_{r}=1$. The bottom-right corner of each image displays the absolute difference between the reconstructed image and the reference image $\boldsymbol{x}_{base}$.
@@ V-A Effectiveness of Relay Residual Diffusion
We first investigate the effectiveness of our proposed relay residual diffusion framework. As shown in Table II, by incorporating the proposed relay residual diffusion framework, the method (+RRD) achieves better distortion metrics (PSNR and MS-SSIM) with only 2 or 5 denoising steps compared to the Baseline, which uses 50 denoising steps. The reason behind this is that starting from the compressed latent feature, instead of pure noise, avoids the error accumulation in the initial stage of the denoising process and provides a solid foundation for subsequent detail generation. The slight drop in perceptual metrics (DISTS and LPIPS) is acceptable, as the relay residual diffusion framework requires significantly fewer denoising steps. Moreover, as shown in Fig. 8(b), when using only 2 denoising steps, the baseline method—which starts from pure noise—fails to completely remove the noise, resulting in noticeable artifacts in the reconstructed image. In contrast, our relay residual diffusion framework constructs the starting point from compressed latent features, enabling high-quality reconstruction even with 2 denoising steps, as shown in Fig. 8(c).
@@
It is worth noting that, since the time required for the denoising stage is directly proportional to the number of denoising steps, incorporating the RRD framework reduces the denoising time by a factor of 10$\times$ to 25$\times$ compared to the baseline, as shown in Table II.
V-B Effectiveness of Fixed-step Fine-tuning
We further investigate the effectiveness of the proposed FSFT strategy. Compared to existing diffusion-based extreme image compression methods, our proposed FSFT strategy fine-tunes the model using the entire reconstruction process and directly applies constraints on the final estimated noise-free latent features $\hat{\boldsymbol{z}}_{0}$ and the reconstructed image $\hat{\boldsymbol{x}}$, as shown in Eq. 15. This strategy is aligned with the actual inference process and offers two primary advantages. On the one hand, by training along a reconstruction trajectory consistent with inference, FSFT effectively mitigates the mismatch between the training and inference phases, thereby enhancing the model’s stability and performance in practical scenarios. On the other hand, by directly constraining the latent features and reconstructed image, rather than relying on noise estimation errors, FSFT enables the model to explicitly learn the degradation patterns inherent in compressed latent features, facilitating more accurate recovery of lost image details.
As demonstrated in Table II, incorporating the FSFT strategy significantly improves performance across all evaluated metrics, confirming that FSFT effectively addresses the training-inference mismatch present in existing diffusion-based compression methods. Furthermore, as shown in Fig. 8(d), the model fine-tuned with the FSFT strategy demonstrates an enhanced ability to recover fine-grained image details, such as clothing folds. It is worth noting that FSFT is a fine-tuning strategy that introduces no additional computational overhead during inference.
Figure 10: Impact of $\lambda_{s}$ on quantitative metrics. All metrics are computed on the CLIC2020 dataset. For the computation of $D$, we adopt the reconstruction result of RDEIC-2 at $\lambda_{s}=0$ (denoted ad $\boldsymbol{x}_{base}$) as the reference image.
(a) Original (b) Full-image (c) Block-wise
[…]
Figure 11: Reconstruction results at different resolutions. (b) the result obtained by directly compressing the entire image, (c) the result obtained by compressing the image patch-by-patch and then stitching the blocks together.
V-C Smoothness-sharpness Trade-off
To provide a more comprehensive analysis of the impact of $\lambda_{s}$, we further present the corresponding performance curves in Fig. 10. As $\lambda_{s}$ increases, both perceptual quality and reconstruction fidelity initially improve but subsequently decline. To further explain this phenomenon, we adopt the reconstruction result of RDEIC-2 at $\lambda_{s}=0$ (denoted as $\boldsymbol{x}_{base}$) as the reference image, and compute a distance metric $D$ to quantify the amount of high-frequency content introduced into the reconstructed image $\boldsymbol{x}_{\lambda_{s}}$:
$D=\|\boldsymbol{x}_{\lambda_{s}}-\boldsymbol{x}_{base}\|_{2}^{2}.$ (18)
As shown in Fig. 10, the value of $D$ increases monotonically with $\lambda_{s}$, indicating that more high-frequency information is introduced into the reconstructed image as $\lambda_{s}$ grows. Initially, introducing a moderate amount of high-frequency detail enhances structural expressiveness and visual realism, resulting in improved perceptual quality and fidelity metrics. Subsequently, when $\lambda_{s}$ becomes excessively large, the generated high-frequency content may deviate from the original image, resulting in mismatched textures and a subsequent decline in both perceptual quality and distortion performance. Furthermore, the distance curve of RDEIC-5 is flatter than that of RDEIC-2, suggesting that increasing the number of denoising steps improves the model’s robustness to variations in conditional intensity.
V-D Robustness to Different Image Resolutions
To validate the robustness of our RDEIC across different image resolutions, we conduct experiments on image patches with varying sizes and aspect ratios. Fig. 11(b) presents the result obtained by directly compressing the entire image, while Fig. 11(c) shows the result obtained by compressing the image patch-by-patch and then stitching the blocks together. These visual results demonstrate that our method effectively supports image compression across various dimensions and resolutions, such as $256\times 256$, $512\times 768$, and $896\times 1408$, while maintaining consistent visual quality and compression performance.
@@ V-D Robustness to Different Image Resolutions
Original W/o diffusion mechanism W/o denoising process RDEIC-5
@@ VI Conclusion
In this paper, we propose an innovative relay residual diffusion-based method (RDEIC) for extreme image compression. Unlike most existing diffusion-based methods that start from pure noise, RDEIC initializes with the compressed latent features of the input image, supplemented with added noise, and reconstructs the image by iteratively removing the noise and reducing the residual between the compressed latent features and the target latent features. To address the discrepancy between the training and inference phases, we propose to fine-tune the model using the entire reconstruction process, further enhancing performance. Inspired by classifier-free guidance (CFG), we introduce a method to balance smoothness and sharpness, overcoming the fixed-step constraint introduced by FSFT. This approach empowers users to explore and customize outputs according to their personal preferences. Extensive experiments have demonstrated the superior performance of RDEIC compared to state-of-the-art methods in terms of both reconstruction quality and computational efficiency.